1/2

Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева ,

Ode a Samarcanda Ode to Samarkand  Oда Самарканду
tradukita de Carlo Minnaja tradukita de Malcolm Jones

Argiope-informo: La aspekto de tiu ĉi versio

estas barita. Supozeble la aŭtoro decidis

tion.

I sing of you, oh, Samarkand’! Your wondrous features win my

praises. And your dear heart - the Registan

- seems young to me and never ages.

Пою тебя, о, Самарканд! Твой чудный облик

восхваляю. И твоё сердце - Регистан - Я вечно юным называю.

Its steady beating Ulughbek could sense beneath the constella-

tions. Oh, Samarkand! Immortal pick of mighty Tamerlan’s creations.

Его биенье Улугбек

Предчувствовал под небом

звёздным, О, Самарканд! Твой

долгий век Величием Тимура создан!

Madrasses are your sacred schools, which cultivate the arts and lear-

ning. Their beauty overcomes time’s ru-

les by centuries into moments turning.

Твои святые медресе - Приют искусств и

просвещенья, Они стоят во всей красе Столетья превратив в

мгновенья.

Here the great Navoi’s verse, so full of love and strong emotion, is cherished by your populace, made happy by its inspiration.

Здесь Алишера Навои Поэзию - любовь и силу - Впитали жители твои Она их счастьем

вдохновила.

The domes of minarets stand proud

- like day-time stars in clear sky shi-

ning: so pure the blue, without a cloud and fine the spires in art combi-

ning.

Здесь минаретов купола - Дневные звёзды в небе

ясном - Так чистая лазурь светла И вязь орнаментов

прекрасна.

…  …  …

2/2

Oh, Samarkand! You’re wise

through age, with mosques and churches closely

neighboured. For those who love you, the world’s

stage has no town better nor more favou-

red.

О, Самарканд! Ты - мудрость лет,

Здесь рядом церкви и

мечети… В тебя влюблённым лучше

нет И краше города на свете!

Oh, Samarkand! Live long and

thrive, be glorified by each good action! And while eternity’s your prize, be here our lasting benefaction.

О, Самарканд! Живи,

цвети И славься добрыми

делами! Навстречу вечности лети И вечно оставайся с нами.

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en itala de Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.- сост.: А. Ионесов . – Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_Minnaja .

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Anglan de Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEA-jarlibro de la jaro 2014 sur paĝo 126.

Verkinto de tiu ĉi Rusa poemo estas Ирина Саверьевна Алексеева .

Arg-860-1733 (2014-05-15 12:22:00)

Tiu ĉi poemo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов .

– Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил.

http://www.poezio.net/version?poem-id=860&version-id=2206,2037,1733